andrea and alexis are twins. they have a brother who is eleven years older than they are and a brother who is four years younger. the sum of the ages of all four children is 35. write an equation to finds sisters age.
Answer by Cromlix(4381)

Andrea and Alexis age = x
Older brother = 11 + x
Younger brother = x - 4)
Total: x + x + (11 + x) + (x - 4) = 35
x + x + 11 + x + x - 4 = 35
4x + 7 = 35
4x = 28
x = 7
Sisters' ages = 7 years each
Younger brother = 3 years
Older brother = 18 years
Hope this helps.
:-)